### v1

@@ 3.2 Asymptotic null distributions
Suppose Assumption 1, 7 and 8 hold. Then, under the null hypothesis in (6),
@@
where $S_{\infty}(\cdot,\cdot)$ is a zero mean Gaussian process with covariance kernel $\mathbb{E}\left[S_{\infty}(\mu,\nu)S_{\infty}(\mu^{\prime},\nu^{\prime})\right]$ and $\rightsquigarrow"$ denotes the weak convergence.
@@
Suppose Assumption 1, 7 and 8 hold. Then, under the null hypothesis in (6),
@@ 3.3 Consistency and asymptotic local power
Suppose Assumption 1, 7 and 8 hold. Then, under the alternative hypothesis, for each $(\mu,\nu)\in\mathcal{W}$,
@@
Under the alternative, since $\mathbb{E}\left[\left(Y_{t}-m(\boldsymbol{Y}_{t-1})\right)e^{iw^{\top}\boldsymbol{W}_{t-1}}\right]$ in a set with positive Lebesgue measure, our statistic $\mathrm{KS}_{n}$ diverges to infinity and achieves asymptotic power of one.
@@
$\mathrm{H}_{1n}:Y_{t}=m(\boldsymbol{Y}_{t-1})+n^{-1/2}\Delta(\boldsymbol{W}_{t-1})+\varepsilon_{t},\quad\mathrm{a.s.}$ (11)
where $\mathbb{E}(\varepsilon_{t}|\boldsymbol{W}_{t-1})=0$, and $\Delta$ is a non-constant measurable function, satisfying $Q(\mu,\nu)=\mathbb{E}(\Delta(\boldsymbol{W}_{t-1})e^{iw^{\top}\boldsymbol{W}_{t-1}})\neq 0$ a.s. for $w\in\mathcal{W}$. Recall that $w=(\mu,\nu)$. The form of local alternatives in (11) is commonly used when studying the asymptotic local power of a test based on empirical processes.
@@
Suppose Assumption 1, 7, 8 and 9 hold. Then, under the sequence of local alternatives in (11),
@@ 3.4 Bootstrap
where $\left\{\xi_{t}\right\}_{t=1}^{n}$ is a sequence of i.i.d. random variables with zero mean, unit variance, and bounded support, independent of $\left\{\left(X_{t},Y_{t}\right)^{\top}\right\}_{t=1}^{n}$.
@@
Theorem 4 implies that the limiting behavior of $\widehat{S}_{n}(\mu,\nu)$ can be approximated by $\widehat{S}^{*}_{n}(\mu,\nu)$. Consequently, the bootstrap-assisted test statistic $\mathrm{KS}_{n}^{*}$ has the same convergence of $\mathrm{KS}_{n}$, constructed by substituting $\widehat{S}_{n}^{*}(\mu,\nu)$ for $\widehat{S}_{n}(\mu,\nu)$ in (10). Therefore, the test statistic $\mathrm{KS}_{n}^{*}$ achieves the correct asymptotic significance level. In addition, $\mathrm{KS}_{n}^{*}$ is consistent against the alternative, and is able to detect local alternatives in (11). The detailed bootstrap procedure is already presented in Algorithm 1.
@@ 4 Simulation
Henceforth, the terms lag order and lag length are used interchangeably and denoted by $L_{a}$. The parameter $L_{a}$ specifies the number of lags of $Y_{t}$ or $X_{t}$, and thus defines the dimensionality of the input to the deep learning model, and in our setting it satisfies that $L_{a}=p=q$. For each DGP, we examine five scenarios corresponding to $L_{a}=1,2,3,4$ and $5$. The following is the specification of the parameters, and we summarize the settings in Table 2. For $L_{a}=1$, we set $a_{1}=1$, $b_{1}=-1$, $c_{1}=-1$, $a=0.5$, $c=1$, and $a_{0}=0.5$. For $L_{a}=2$, we set $(a_{1},a_{2})^{\top}=(0.5,-0.5)^{\top}$, $(b_{1},b_{2})^{\top}=(-0.5,0.5)^{\top}$, $(c_{1},c_{2})^{\top}=(-0.5,0.5)^{\top}$, $a=0.25$, $c=0.6$, and $a_{0}=0.4$. For $L_{a}=3$, we set $(a_{1},a_{2},a_{3})^{\top}=(0.5,-0.5,0.5)^{\top}$, $(b_{1},b_{2},b_{3})^{\top}=(-0.5,0.5,0.5)^{\top}$, $(c_{1},c_{2},c_{3})^{\top}=(-0.5,0.5,-0.5)^{\top}$, $a=0.25$, $c=0.5$, and $a_{0}=1/3$. For $L_{a}=4$, we set $(a_{1},a_{2},a_{3},a_{4})^{\top}=(0.25,-0.25,0.25,0.25)^{\top}$, $(b_{1},b_{2},b_{3},b_{4})^{\top}=(-0.25,0.25,0.25,-0.25)^{\top}$, $(c_{1},c_{2},c_{3},c_{4})^{\top}=(-0.25,0.25,-0.25,0.25)^{\top}$, $a=0.125$, $c=0.5$, and $a_{0}=1/3$. For $L_{a}=5$, we set $(a_{1},a_{2},a_{3},a_{4},a_{5})^{\top}=(0.25,-0.25,0.25,0.25,-0.25)^{\top}$, $(b_{1},b_{2},b_{3},b_{4},b_{5})^{\top}=(-0.25,0.25,0.25,-0.25,0.25)^{\top}$, $(c_{1},c_{2},c_{3},c_{4},c_{5})^{\top}=(-0.25,0.25,-0.25,0.25,-0.25)^{\top}$, $a=0.125$, $c=0.5$, and $a_{0}=1/3$. These parameters are designed to constrain the summation in $Y_{t}$ across multiple lags to prevent it from diverging during data generation, although they may not produce the best performance for our test.
@@
For hyperparameters, we select the number of mixture components $G$ via cross-validation, as it affects empirical performance to a certain extent. A small value of $G$ can introduce a large bias in the fitted MDN model, leading to inflated type I error rates. In contrast, a large $G$ can cause overfitting, resulting in increased variability of the test statistic. In general, the setting $G=10$ is suitable for most scenarios. Regarding the number of pairs $L$, a higher value of $L$ generally increases the power of the test, but at the cost of higher computational complexity. To balance power performance and efficiency, we fix $L=20$. For the number of pseudo-samples $M$, we find that the test is relatively insensitive to this parameter, provided $M$ within a reasonable range. Hence, we fix $M=20$ for all numerical studies.
For both MLP and MDN models, we fix the number of layers $H=1$. The number of nodes $U$ per hidden layer is varied as $U=5L_{a}$. Recall that $L_{a}$ denotes the lag parameter. This choice provides a balance between network simplicity and simulation performance. For the MLP, the loss function can be chosen as $L_{2}$ loss or smooth $L_{1}$ loss, while the parameters are estimated using maximum likelihood for the MDN.
Let $T_{0}=500$. For each DGP, we generate $T_{0}+T$ observations and discard the initial $T_{0}$ observations to mitigate the effects of the initial values. The simulations are based on $1000$ Monte Carlo replications, and the bootstrap $p$-values are computed using $B=1000$ bootstrap replications. Three sample sizes are considered, $T=500$, 1000, and 2000. We report results for a nominal significance level of $5\%$, with results for other levels available upon request.
@@
It is worth noting that the results of size are reasonable when the sample sizes are increasing, such as $T=1000$ and $T=2000$. However, if hastily integrate deep neural network estimators —specifically, by using the hypothesis in (3) to construct a Granger causality test —this approach fails to control the type I error, especially when the sample size increases. For example, in DGP S1, the size is 0.151 for $T=1000$ and 0.321 for $T=2000$ under $L_{a}=5$. This phenomenon highlights the necessity for the doubly robust Granger causality test, which has facilitated the practical integration of deep learning into statistical testing methods. Furthermore, as observed, dimension does not significantly affect the size property under the proposed test. We also employ the test proposed by Nishiyama et al. (2011). Since the dimensionality is up to 5, we adopt a fourth-order Gaussian kernel function for NHKJ to align with the theoretical analysis. To ensure fairness, we set different bandwidths in S1 and S2 across varying lags for NHKJ. For DGP S1, set bandwidth $h=2.5\times T^{-0.15}$ when $L_{a}=1$, and $h=3\times T^{-0.15}$ when $L_{a}=2,3$, while $h=3.5\times T^{-0.15}$ when $L_{a}=4,5$. Similarly, for DGP S2, set bandwidth $h=3\times T^{-0.15}$ when $L_{a}=1$, and $h=3.5\times T^{-0.15}$ when $L_{a}=2,3$, while $h=4\times T^{-0.15}$ when $L_{a}=4,5$. According to the results in Table 3, NHKJ struggles to control type I error at the nominal level and exhibits a trend of increasing type I error as the sample size grows. This may be attributed to NHKJ’s sensitivity to bandwidth selection. We do not choose a sufficiently large bandwidth to control the type I error at the nominal significance level because doing so would require an excessively large value. For example, in the case of S2 with $L_{a}=5$, the bandwidth would need to exceed $8\times T^{-0.15}$.
@@
Then, we investigate the power performance of the proposed test. Specifically, DGP P1, P2, P3 and P4 are designed to evaluate the power while the lag parameter $L_{a}$ varies from 1 to 5, examining the impact of increasing lag orders on power performance. DGP P1, P2, and P3 consider an additive structure for $Y_{t}$, where $Y_{t}$ is modeled as the sum of a pure $\boldsymbol{Y}_{t-1}$ component and a pure $\boldsymbol{X}_{t-1}$ component. Particularly, P1 and P2 are designed with a linear combination of $\boldsymbol{Y}_{t-1}$, while P3 introduces a more complex nonlinear structure for $\boldsymbol{Y}_{t-1}$. Regarding the $\boldsymbol{X}_{t-1}$ component in $Y_{t}$, P1 incorporates a linear summation of $\boldsymbol{X}_{t-1}$, which is then transformed by a nonlinear function, that is $\sin$. In contrast, P2 and P3 apply a nonlinear transformation to $\{X_{t-1},\cdots,X_{t-p}\}$ individually before summation. DGP P4, unlike the other three settings, adopts a more hybrid structure. It includes the product of $\{X_{t-i},Y_{t-i}\}_{i=1}^{q}$, rather than treating $\boldsymbol{X}_{t-1}$ and $\boldsymbol{Y}_{t-1}$ as separate components.
@@
As observed, our method performs strongly when the dimensionality ranges from one to three, with power approaching one. While performance declines with increasing dimensionality, the test remains effective in cases with lag orders of four and five, which are generally considered difficult for nonparametric methods. This indicates a potential improvement in large-lag Granger causality testing. For NHKJ, note that DGP P1 and P2 correspond to DGP S1, while DGP P3 corresponds to DGP S2. Therefore, the bandwidth settings are aligned with those used in S1 and S2, respectively. As shown, the power of NHKJ is significantly affected as the lag order increases—even at $L_{a}=2$, it exhibits noticeably lower power compared to the doubly robust test. This underscores the advantage of our proposed method in testing Granger causality across multiple lags.
@@ 5 Real data application
In this section, we investigate whether Granger causality exists between stock prices and trading volumes using practical data. It is widely recognized in financial markets that there is a relationship between stock price fluctuations and trading volume in adjacent periods —a phenomenon that has been extensively documented in economic and financial research. This price-volume relationship serves as a critical indicator for investors to anticipate future trends. Roughly speaking, trading volume tends to increase during periods of noticeable price changes and decrease when price movements are less pronounced. Additionally, significant price fluctuation often follows periods of high trading volume.
We use daily price and volume data from three renowned stock market indices representing different economies: the S&P 500 Index (SPX 500), the CSI 300 Index (CSI 300), and the Nikkei 225 Index (NI 225). The sample period spans five years, from September 27, 2019 to September 26, 2024, yielding a total sample size of $T=1257$. To ensure stationarity, both price and volume are transformed into percentage changes. For volume, the percentage change is divided by 10 to conform model training and achieve a scale comparable to that of stock prices. We denote these percentage-change variables as $P_{t}$ for price and $V_{t}$ for volume.
To investigate the stability and consistency of Granger causality patterns, we conduct analyses using overlapping three-year sub-samples for each index. Specifically, we examine Granger causality using data from September 27, 2019 to September 26, 2022, from September 27, 2020 to September 26, 2023, and from September 27, 2021 to September 26, 2024, hereafter referred to as 2019-2022, 2020-2023, and 2021-2024, respectively. The sample size for each three-year period is approximately $T=750$.
The primary objective is to investigate whether there exists causality in the mean between changes in trading volume and stock price. Specifically, we aim to detect bidirectional Granger causality at lag orders ranging from 1 to 10 using the proposed test. To test whether price changes Granger-cause volume changes, we treat $P_{t}$ as $X_{t}$ and $V_{t}$ as $Y_{t}$. Conversely, testing whether volume changes Granger-cause price changes reverses these roles. The parameter settings are mostly consistent with those in the simulation studies, except for setting the lag length from 1 to 10, as the computational cost in this application is significantly lower than in the simulations. The upper 5% critical value is employed as the threshold for hypothesis testing.
@@
The numerical results yield several noteworthy conclusions. For each index, we observe a consistent—or at least shifting—pattern of Granger causality in mean from stock prices to trading volumes. This causality does not necessarily arise immediately but often manifests across multiple lag orders. Our proposed method is well suited for detecting such patterns, particularly when causality extends over several lags. A further advantage of DRGCT is its robustness in practical applications: in real-world data the true lag order is typically unknown, making it difficult to determine the appropriate lag length. By employing our test, researchers can set the lag order to a sufficiently large value, thereby covering all potential causal lags.
These findings may also reflect common market behavior. A temporary price change may or may not have an immediate impact on trading activity. By contrast, sustained trends in prices are more likely to influence trading, as investors may base their decisions on cumulative patterns rather than single-day fluctuations.
In the reverse direction, trading volume occasionally serves as a causal factor for stock prices, though such relationships appear far less consistent. It is worth noting that our empirical analysis is restricted to daily data with lag orders up to 10. Thus, causal relationships at longer horizons cannot be ruled out. Moreover, our framework considers only cases where both series share the same lag order, suggesting that more nuanced causal structures may exist and warrant further exploration.

### v2

@@ 3.2 Asymptotic null distributions
Suppose Assumptions 1, 7 and 8 hold. Then, under the null hypothesis in (6),
@@
where $S_{\infty}(\cdot,\cdot)$ is a zero mean Gaussian process with covariance kernel $\mathbb{E}\left[S_{\infty}(\mu,\nu)S_{\infty}(\mu^{\prime},\nu^{\prime})\right]$ and $\rightsquigarrow$” denotes the weak convergence.
@@
Suppose Assumptions 1, 7 and 8 hold. Then, under the null hypothesis in (6),
@@ 3.3 Consistency and asymptotic local power
Suppose Assumptions 1, 7 and 8 hold. Then, under the alternative hypothesis, for each $(\mu,\nu)\in\mathcal{W}$,
@@
Under the alternative, since $\mathbb{E}\left[\left(Y_{t}-m(\boldsymbol{Y}_{t-1})\right)e^{iw^{\top}\boldsymbol{W}_{t-1}}\right]\neq 0$, our statistic $\mathrm{KS}_{n}$ diverges to infinity and achieves asymptotic power of one.
@@
$\mathrm{H}_{1n}:\mathbb{E}\left[Y_{t}-m(\boldsymbol{Y}_{t-1})|\boldsymbol{W}_{t-1}\right]=n^{-1/2}\Delta(\boldsymbol{W}_{t-1}),\quad\mathrm{a.s.}$ (11)
where $\Delta$ is a non-constant measurable function, satisfying $Q(\mu,\nu)=\mathbb{E}(\Delta(\boldsymbol{W}_{t-1})e^{iw^{\top}\boldsymbol{W}_{t-1}})\neq 0$ a.s. for $w\in\mathcal{W}$. Recall that $w=(\mu,\nu)$. The form of local alternatives in (11) is commonly used when studying the asymptotic local power of a test based on empirical processes.
@@ 3.4 Bootstrap
where $\left\{\xi_{t}\right\}_{t=1}^{n}$ is a sequence of i.i.d. random variables with zero mean, unit variance, and bounded support, independent of $\left\{\left(X_{t},Y_{t}\right)^{\top}\right\}_{t=1}^{n}$. Note that if the residual $\left\{Y_{t}-\mathbb{E}(Y_{t}|\boldsymbol{W}_{t-1})\right\}$ is a martingale difference sequence, the choice of $\xi_{t}$ can fully characterize the covariance kernel of the process.
@@
Theorem 4 implies that the limiting behavior of $\widehat{S}_{n}(\mu,\nu)$ can be approximated by $\widehat{S}^{*}_{n}(\mu,\nu)$. Consequently, the bootstrap-assisted test statistic $\mathrm{KS}_{n}^{*}$ has the same convergence of $\mathrm{KS}_{n}$, constructed by substituting $\widehat{S}_{n}^{*}(\mu,\nu)$ for $\widehat{S}_{n}(\mu,\nu)$ in (10). Therefore, given the significance level $\alpha$, the $p$-value of the test can be obtained by calculating the test statistic $\mathrm{KS}_{n}^{*}$ according to Step 4 in Algorithm 1. In addition, $\mathrm{KS}_{n}^{*}$ is consistent against the alternative, and is able to detect local alternatives in (11). See the detailed testing procedure in Algorithm 1.
@@ 4 Simulation
Henceforth, the terms lag order and lag length are used interchangeably and referred to simply as $\mathrm{lag}$. $\mathrm{Lag}$ specifies the number of lags of $Y_{t}$ or $X_{t}$, and thus defines the dimensionality of the input to the deep learning model, and in our setting it satisfies that $\mathrm{lag}=p=q$. For each DGP, we examine five scenarios corresponding to $\mathrm{lag}=1,2,3,4$ and $5$. The following is the specification of the parameters, and we summarize the settings in Table 2. For $\mathrm{lag}=1$, we set $a_{1}=1$, $b_{1}=-1$, $c_{1}=-1$, $a=0.5$, $c=1$, and $a_{0}=0.5$. For $\mathrm{lag}=2$, we set $(a_{1},a_{2})^{\top}=(0.5,-0.5)^{\top}$, $(b_{1},b_{2})^{\top}=(-0.5,0.5)^{\top}$, $(c_{1},c_{2})^{\top}=(-0.5,0.5)^{\top}$, $a=0.25$, $c=0.6$, and $a_{0}=0.4$. For $\mathrm{lag}=3$, we set $(a_{1},a_{2},a_{3})^{\top}=(0.5,-0.5,0.5)^{\top}$, $(b_{1},b_{2},b_{3})^{\top}=(-0.5,0.5,0.5)^{\top}$, $(c_{1},c_{2},c_{3})^{\top}=(-0.5,0.5,-0.5)^{\top}$, $a=0.25$, $c=0.5$, and $a_{0}=1/3$. For $\mathrm{lag}=4$, we set $(a_{1},a_{2},a_{3},a_{4})^{\top}=(0.25,-0.25,0.25,0.25)^{\top}$, $(b_{1},b_{2},b_{3},b_{4})^{\top}=(-0.25,0.25,0.25,-0.25)^{\top}$, $(c_{1},c_{2},c_{3},c_{4})^{\top}=(-0.25,0.25,-0.25,0.25)^{\top}$, $a=0.125$, $c=0.5$, and $a_{0}=1/3$. For $\mathrm{lag}=5$, we set $(a_{1},a_{2},a_{3},a_{4},a_{5})^{\top}=(0.25,-0.25,0.25,0.25,-0.25)^{\top}$, $(b_{1},b_{2},b_{3},b_{4},b_{5})^{\top}=(-0.25,0.25,0.25,-0.25,0.25)^{\top}$, $(c_{1},c_{2},c_{3},c_{4},c_{5})^{\top}=(-0.25,0.25,-0.25,0.25,-0.25)^{\top}$, $a=0.125$, $c=0.5$, and $a_{0}=1/3$. These parameters are designed to constrain the summation in $Y_{t}$ across multiple lags to prevent it from diverging during data generation, although they may not produce the best performance for our test.
@@
For hyperparameters, we select the number of mixture components $G$ via cross-validation, as it affects empirical performance to a certain extent. A small value of $G$ can introduce a large bias in the fitted MDN model, leading to inflated type I error rates. In contrast, a large $G$ can cause overfitting, resulting in increased variability of the test statistic. In general, the setting $G=10$ is suitable for most scenarios. Regarding $L$, that is the number of $(\mu,\nu)$ pairs, a higher value of $L$ generally increases the power of the test, but at the cost of higher computational complexity. To balance power performance and efficiency, we fix $L=20$. For the number of pseudo-samples $M$, we find that the test is relatively insensitive to this parameter, provided $M$ within a reasonable range. Hence, we fix $M=20$ for all numerical studies.
For both MLP and MDN models, we fix the depth of neural networks $L_{n}$ to be $L_{n}=1$. The width, that is the number of nodes $H_{n}$ per hidden layer, is varied with the lag as $H_{n}=5\cdot\mathrm{lag}$. This choice provides a balance between network simplicity and simulation performance. For the MLP, the loss function can be chosen as $L_{2}$-loss or smooth $L_{1}$-loss, while the parameters are estimated using maximum likelihood for the MDN.
Let $n_{0}=500$. For each DGP, we generate $n_{0}+n$ observations and discard the initial $n_{0}$ observations to mitigate the effects of the initial values. The simulations are based on $1000$ Monte Carlo replications, and the bootstrap $p$-values are computed using $B=1000$ bootstrap replications. Three sample sizes are considered, $n=500$, 1000, and 2000. We report results for a nominal significance level of $5\%$, with results for other levels available upon request.
@@
It is worth noting that the results of size are reasonable when the sample sizes are increasing, such as $n=1000$ and $n=2000$. However, if hastily integrate deep neural network estimators —specifically, by using the hypothesis in (3) to construct a Granger causality test —this approach fails to control the type I error, especially when the sample size increases. For example, in DGP S1, the size is 0.151 for $n=1000$ and 0.321 for $n=2000$ under $\mathrm{lag}=5$. This phenomenon highlights the necessity for the doubly robust Granger causality test, which has facilitated the practical integration of deep learning into statistical testing methods. Furthermore, as observed, dimension does not significantly affect the size property under the proposed test. We also employ the test proposed by Nishiyama et al. (2011). Since the dimensionality is up to 5, we adopt a fourth-order Gaussian kernel function for NHKJ to align with the theoretical analysis. To ensure fairness, we set different bandwidths in S1 and S2 across varying lags for NHKJ. For DGP S1, set bandwidth $h=2.5\times n^{-0.15}$ when $\mathrm{lag}=1$, and $h=3\times n^{-0.15}$ when $\mathrm{lag}=2,3$, while $h=3.5\times n^{-0.15}$ when $\mathrm{lag}=4,5$. Similarly, for DGP S2, set bandwidth $h=3\times T^{-0.15}$ when $\mathrm{lag}=1$, and $h=3.5\times n^{-0.15}$ when $\mathrm{lag}=2,3$, while $h=4\times n^{-0.15}$ when $\mathrm{lag}=4,5$. According to the results in Table 3, NHKJ struggles to control type I error at the nominal level and exhibits a trend of increasing type I error as the sample size grows. This may be attributed to NHKJ’s sensitivity to bandwidth selection. We do not choose a sufficiently large bandwidth to control the type I error at the nominal significance level because doing so would require an excessively large value. For example, in the case of S1 with $\mathrm{lag}=5$, the bandwidth should be set to $4\times n^{-0.15}$, whereas in the case of S2 with $\mathrm{lag}=5$, the bandwidth is required to exceed $8\times n^{-0.15}$.
@@
Then, we investigate the power performance of the proposed test. Specifically, DGP P1, P2, P3 and P4 are designed to evaluate the power while the lag parameter $\mathrm{lag}$ varies from 1 to 5, examining the impact of increasing lag orders on power performance. DGP P1, P2, and P3 consider an additive structure for $Y_{t}$, where $Y_{t}$ is modeled as the sum of a pure $\boldsymbol{Y}_{t-1}$ component and a pure $\boldsymbol{X}_{t-1}$ component. Particularly, P1 and P2 are designed with a linear combination of $\boldsymbol{Y}_{t-1}$, while P3 introduces a more complex nonlinear structure for $\boldsymbol{Y}_{t-1}$. Regarding the $\boldsymbol{X}_{t-1}$ component in $Y_{t}$, P1 incorporates a linear summation of $\boldsymbol{X}_{t-1}$, which is then transformed by a nonlinear function, that is $\sin$. In contrast, P2 and P3 apply a nonlinear transformation to $\{X_{t-1},\cdots,X_{t-p}\}$ individually before summation. DGP P4, unlike the other three settings, adopts a more hybrid structure. It includes the product of $\{X_{t-i},Y_{t-i}\}_{i=1}^{q}$, rather than treating $\boldsymbol{X}_{t-1}$ and $\boldsymbol{Y}_{t-1}$ as separate components.
@@
As observed, our method performs strongly when the dimensionality ranges from one to three, with power approaching one. While performance declines with increasing dimensionality, the test remains effective in cases with lag orders of four and five, which are generally considered difficult for nonparametric methods. This indicates a potential improvement in large-lag Granger causality testing. For NHKJ, note that DGP P1 and P2 correspond to DGP S1, while DGP P3 corresponds to DGP S2. Therefore, the bandwidth settings are aligned with those used in S1 and S2, respectively. As shown, the power of NHKJ is significantly affected as the lag order increases—even at $\mathrm{lag}=2$, it exhibits noticeably lower power compared to the doubly robust test. This underscores the advantage of our proposed method in testing Granger causality across multiple lags.
@@ 5 Real data application
In this section, we examine the presence of Granger causality between stock prices and trading volumes using empirical data. It is well established in the finance literature that stock price fluctuations and trading volumes are closely related across adjacent periods —a phenomenon extensively documented in both economics and financial research (see Karpoff, 1987; Gallant et al., 1992; Chen et al., 2001). This price–volume relationship is widely regarded as an important indicator for anticipating future market trends. In general, trading volumes tend to rise during periods of substantial price movements and decline when price changes are less pronounced. Conversely, periods of high trading volume are often followed by significant price fluctuations.
We use daily prices and volumes data from three renowned stock market indices representing different economies: the S&P 500 Index (SPX 500), the CSI 300 Index (CSI 300), and the Nikkei 225 Index (NI 225). The sample period spans five years, from September 27, 2019 to September 26, 2024, yielding a total sample size of $T=1257$. To ensure stationarity, both prices and volumes are transformed into percentage changes. For volumes, the percentage change is divided by 10 to conform model training and achieve a scale comparable to that of stock prices. We denote these percentage-change variables as $P_{t}$ for prices and $V_{t}$ for volumes.
To investigate the stability and consistency of Granger causality patterns, we conduct analyses using overlapping three-year sub-samples for each index. Specifically, we examine Granger causality using data from September 27, 2019 to September 26, 2022, from September 27, 2020 to September 26, 2023, and from September 27, 2021 to September 26, 2024, hereafter referred to as 2019-2022, 2020-2023, and 2021-2024, respectively. The sample size for each three-year period is approximately $n=750$.
The primary objective is to investigate whether there exists causality in the mean between changes in trading volumes and stock prices. Specifically, we aim to detect bidirectional Granger causality at lag orders ranging from 1 to 10 using the proposed test. To test whether price changes Granger-cause volume changes, we treat $P_{t}$ as $X_{t}$ and $V_{t}$ as $Y_{t}$. Conversely, testing whether volume changes Granger-cause price changes reverses these roles. The parameter settings are mostly consistent with those in the simulation studies, except for setting the lag length from 1 to 10, as the computational cost in this application is significantly lower than in the simulations. The upper 5% critical value is employed as the threshold for hypothesis testing.
@@
The numerical results yield several noteworthy conclusions. For each index, we observe either a consistent or a shifting pattern of Granger causality in mean from stock prices to trading volumes. The characteristics differ across indices: for the SPX 500, there appears to be persistent causality from prices to volumes; for the CSI 300, the Granger causality appears to evolve over time, becoming more pronounced in later periods. and for the Nikkei 225, no evidence of causality is found across all periods. Notably, such causality does not necessarily emerge immediately but often manifests over multiple lag orders. Our proposed method is particularly effective in capturing these patterns, especially when causality extends across several lags. An additional advantage of DRGCT is its robustness in practical applications: since the true lag order is typically unknown in real-world data, selecting the appropriate lag length can be challenging. By applying our test, researchers can set the lag order to a sufficiently large value, thereby encompassing all potential causal lags.
These findings may also reflect general market behavior. A temporary price change does not necessarily exert an immediate impact on trading activity. By contrast, sustained price trends are more likely to affect trading, as investors tend to base their decisions on cumulative patterns rather than isolated daily fluctuations. This interpretation is consistent with Copeland (1976), whose sequential information arrival model shows that trading volume responds more strongly to the cumulative impact of information arrivals rather than to isolated shocks, and also relates to De Long et al. (1990), who demonstrate how noise traders can cause persistent price deviations and market anomalies.
In the reverse direction, trading volumes occasionally serve as a causal factor for stock prices, though such relationships appear far less consistent. It is worth noting that our empirical analysis is restricted to daily data with lag orders up to 10. Thus, causal relationships at longer horizons cannot be ruled out. Moreover, our framework considers only cases where both series share the same lag order, suggesting that more nuanced causal structures may exist and warrant further exploration.